OpenStudy (anonymous):

Rationalize the denominator. The answer should be expressed in simplified form.

OpenStudy (anonymous):

\[\frac{ \sqrt{5}-\sqrt{7} }{ \sqrt{5}+\sqrt{3}}\]

Miracrown (miracrown):

@cookiibabii93 I assume you've done many similar problems like this; just a couple of mins ago we were working on a similar one. Do you think you would be able to handle this on your own now? - or are you still seeking for help?

OpenStudy (anonymous):

No, it's only 2

OpenStudy (anonymous):

Step by step

Miracrown (miracrown):

Oh lol ok :-)

OpenStudy (anonymous):

I know I multiply everything by the denominator

OpenStudy (paki):